# ElastiBench: Scalable Continuous Benchmarking on Cloud FaaS Platforms

### v1

@@ 6.1. Experiment Design
Our evaluation focuses on the open-source time-series database software VictoriaMetrics, as we have used it in previous work and have reliable performance results to compare ElastiBench to (Grambow et al., 2023a). We repeat the same microbenchmarks with the same SUT versions that were used in the original paper and compare their results.
@@
For this evaluation, all microbenchmarks which collect less than 10 results are ignored. Microbenchmarks might fail because they cannot run (e.g., a build failure), or because they ran for more than twenty seconds, after which they are interrupted. This can happen for some benchmarks that require extensive setups or have long iterations. We argue that these are not relevant as single microbenchmark executions microbenchmarks longer than one second with the default Go benchmark parameters are indicative of faulty benchmark implementations (Cheney, 2019). In an experiment, multiple microbenchmarks are run. Some microbenchmark executions are run for different configurations, e.g., different input sizes for a function. We treat these configurations as independent microbenchmarks. If the 99% confidence interval (CI) of the median difference in performance between the two SUT versions that we calculated using bootstrapping (see Section 2) does not overlap 0, then the experiment has detected a performance change for that microbenchmark. If one experiment detects a performance change and another one does not, we call this a possible performance change, as it was not reliably detected across multiple experiments. Figure 3 (note the x-axis values) shows an example of a distribution that has a performance change as opposed to no performance change.
Figure 3. When comparing the difference in performance between two versions, the confidence interval shows range where the real difference likely is. If the CI overlaps zero (black curve), we detect no performance change. If it does not overlap zero (yellow curve), we detect a performance change
@@
Unless otherwise noted, we deploy all functions with a timeout of 15151515min, which is the maximum timeout allowed by AWS Lambda. The functions have 2048 MB RAM, and access to 1.29 vCPUs using an ARM architecture. Every function call contains one microbenchmark call, which is repeated three times inside the function. This call is repeated 15 times per microbenchmark, leading to 45 results for every microbenchmark. This is the same number used in our previous study (Grambow et al., 2023a) and above the number of repeats found in other research (Laaber et al., 2019). We limit the call parallelism of the call script to 150, i.e., a maximum of 150 microbenchmarks execute on the FaaS platform in parallel.
Every microbenchmark is repeated until a stable number of executions per second is found. The SUT in these experiments is the open-source time series database VictoriaMetrics. The two software versions we compare are the initial (f611434) and last commit (7ecaa2fe) evaluated in our previous work (Grambow et al., 2023a), which we refer to as the original dataset in this paper. In this paper, we run the following experiments:
  1. (1)
    A/A Experiment: In the A/A experiment, both versions of the SUT are the initial commit used in our previous work (Grambow et al., 2023a). With this experiment, we verify that the system works and that the variability of the FaaS platform does not invalidate the results.
  2. (2)
@@ 6.2. Experiment Results
We first report the results of the all experiments that collect the same data as the original dataset (Sections 6.2.1, 6.2.2, 6.2.4, 6.2.3 and 6.2.5), and then compare their results against each other (Section 6.2.6). Afterwards, we analyze how many repeats are necessary to achieve a consistent CI size (Section 6.2.7).
6.2.1. A/A Experiment
@@
Out of 106 microbenchmarks in the SUT, 90 were executed successfully. Out of those, no performance change was found. This shows that ElastiBench does not mistakenly detect performance changes in unchanged software. Figure 4 shows an overview of the performance differences found. The median difference in performance is 0.047%, with a maximum of 32%. Large performance differences indicate that a microbenchmark is highly variable (as the same version is executed twice), which is correctly detected as no performance change. Running the experiment took ~8min and cost \$1.18444It was started ~2024-05-12T17:35:00 UTC.
@@
Figure 4. CDF showing the performance differences identified in the A/A experiment. While some microbenchmarks have a high difference, all of them are correctly categorized as no performance change.
@@
Figure 5. CDF showing the performance differences identified in the baseline experiment. Performance changes have a generally higher difference than non-changes, with the median performance change being 3.08%.
6.2.2. Baseline Experiment
@@
We show an overview of the differences in performance measured in the baseline experiment in Figure 5. The median performance change of the two SUT versions is 4.71%. The maximum difference in performance of a change and no change are 116% and 26% respectively.
@@
After removing microbenchmarks for which only one experiment contains results, there are 91 (out of 106) microbenchmarks that can be compared to the original dataset. Out of those, both datasets agree for 95.65% of microbenchmarks (88 overall). The disagreements are as follows: for 3.2% of microbenchmarks (three overall), both find a performance change but identify different directions. These three results are from a single microbenchmark function (BenchmarkAddMulti) with different configurations. In every case, the original dataset identifies a performance change of ~-10%, while the new experiments identifies a positive 5-7% performance change. This benchmark was changed between the two SUT versions555commit hash eb103e15 with a considerable change in the benchmark itself, which leads to inconsistent results as the two versions are essentially different benchmarks. We argue that this is not representative for microbenchmarks, so that the difference in results between the original dataset and our experiments are not indicative of failure of ElastiBench.
@@
The remaining disagreement is a single microbenchmark for which the original dataset does not find a performance change while the new dataset finds one. The difference in performance is 1.96% in the original dataset, and 0.60% in the new dataset. As previous research has shown that a performance difference this small is not reliably detectable with cloud microbenchmarks (cf. Section 2), we argue that while this is technically a disagreement, it is an issue of cloud benchmarking in general.
@@
The median performance change measured in this paper has a one-sided coverage with the original dataset for 86.96% of microbenchmarks, and 52.17% vice versa. The two-sided coverage is 50% of microbenchmarks, i.e., for 50% of performance changes in a microbenchmark, the result of one experiment is within the CI of the other result. We argue that the overall high agreement of the baseline experiment with the original dataset demonstrates that both find the same performance changes. The magnitude of the performance difference however depends on many factors that were changed between the baseline experiment and our original work(Grambow et al., 2023a), e.g., the execution environment and version of the programming language, which explains the comparatively low two-sided coverage. The experiment took ~11min, and the total cost for the experiment is ~0.18\$, i.e., a similar cost to the original experiments while taking ~4.6% of time [^6] It was started ~2024-05-12T16:50:00 UTC.
6.2.3. Replication Experiment
@@
The replication experiment has the same agreement with the original dataset as the baseline dataset. The one-sided coverage with the original dataset is 81.72%, and 51.61% in the other direction. The two-sided coverage is just 48.39% It disagrees with the baseline experiment for 10.87% of microbenchmarks, which are all microbenchmarks that are not in the original dataset (i.e., they were not run successfully or had too few runs to be included). The maximum difference in performance found by one experiment where the other experiment did not find a performance change (the possible performance change) was 5.25%. Overall, the results show that while the changes with a higher performance difference are detected continuously, smaller changes are not detected robustly. This maximum possible performance change of 5.25% is still within the limits found in other work (cf. Section 2). The experiment cost \$1.18 to execute and took ~9min [^7] It was started ~2024-05-12T19:35:00 UTC.
6.2.4. Lower Memory Experiment
@@
6.2.5. Single Repeat Experiment
@@
6.2.6. Possible Performance Changes
Figure 6. The maximum difference in performance of a microbenchmark when two experiments disagree whether a performance change has happened.
We compare the difference in performance of all microbenchmarks where the baseline, replication, lower memory, and single-repeat experiments disagree with each other and collect the maximum performance change that was found by one experiment. An overview of these differences can be found in Figure 6. The median difference is 1.58%, the 75th percentile is 3.06%, and the biggest difference is 7.6% in the BenchmarkAdd/items_100000 microbenchmark (which is unreliable, cf. Section 6.2.2) between the single-repeat and lower memory experiments.
6.2.7. Repetitions Necessary for Consistent CI Size
Figure 7. If the CI of the original dataset and the experiment overlap (i.e., they show a similar performance), this graph shows how many repetitions are necessary until the size of the CI of ElastiBench is equal to or less than the size the original dataset. The graph increases until 40 repetitions and then plateaus to around 90% with more and more repetitions
To evaluate how many repetitions of a single microbenchmark are necessary to achieve reliable results, we run an experiment where we repeat every microbenchmark 200 times: 4 times within a function call, which is called 50 times. We then calculate the median and its CI repeatedly with more and more of the calls, which leads to increasingly accurate results. Figure 7 shows how close the results of ElastiBench can get to the original dataset. Over all benchmarks where the CIs ultimately overlap each other (i.e., they share at least one common value), we calculate how many repetitions are necessary until the size of the CI is less than or equal to the CI size of the original dataset. In 75.95% of cases, this is achieved when using the same 45 repeats used in the original dataset. At the full 135 steps, it is achieved for 89.87% of microbenchmarks.
6.3. Key Takeaways
Our evaluation of ElastiBench shows that it is viable to execute microbenchmarking experiments on FaaS platforms, decreasing microbenchmark suite execution time to 6% and lowering total execution costs compared to virtual machine-based experiments. The different experiments show that the platform and experiment configuration seem to have minimal impact on performance changes with a sufficiently large difference that they can reliably detected using cloud platforms. For effect sizes of more than 7.06% results stayed consistent between all experiment runs, with a median inconsistent change of 1.63% (cf. Section 6.2.6). Other researchers have found that a relatively small significant performance change between 3% and 10% does not imply a real-world change due to the underlying variability in cloud platforms and microbenchmarks (Georges et al., 2007; Mytkowicz et al., 2009; Laaber et al., 2019). Our results fall within that range.
7. Discussion & Future Work
@@
7.1. Optimizing Function Configuration
@@
We use functions with 2048 MB of memory in our evaluation to ensure no microbenchmark runs out of memory, yet not every microbenchmark necessarily requires these resources. Previous research has shown that it is possible to optimize FaaS function resource allocation further while the application is running (Schirmer et al., 2022). By adapting the required resources per microbenchmark function, the total execution cost of the microbenchmark suite could be reduced further. As most cloud FaaS platforms allocate compute resources with the configured memory size, however, care must be taken to not influence benchmark results, e.g., as using shared CPU cores could further increase performance variability. Overall, however, there are limits to how low we can go in terms of resources as the SUT is, in the end, still a complete system such as VictoriaMetrics in our experiments.
7.2. Benchmarking Strategy
@@
Our current benchmarking strategy with 45 repetitions of each microbenchmark (cf. Section 6.2.7) is sufficient to reduce the mean standard error of results that show a performance change to less than two percent, with an overall achievable standard error of around one percent. In future research, we plan to further refine this strategy in order to increase the cost and time savings of ElastiBench. We plan to evaluate the impact of repeating the microbenchmark within a function instance compared to repeatedly invoking a function. Additionally, we will further assess the impact of cold start performance that results from calling multiple benchmarks in parallel.
7.3. Microbenchmarks Beyond Go
@@
Our current prototype supports Go microbenchmarks, yet Go is not the only programming language with support for microbenchmarks: Future extensions to our prototype could support Rust’s cargo-bench (Rust Contributors, 2024), Benchmark.js (BestieJS Modules, 2024), Java’s Microbenchmark Harness(OpenJDK Contributors, 2024), or the Google Benchmark Framework for C++ projects (Google, 2024). Integrating such microbenchmark frameworks into ElastiBench will require further evaluation of their interaction with cloud FaaS performance variability.
7.4. Limitations of the Lambda Runtime
@@
As described in Section 3, the execution environment of cloud FaaS platforms is typically locked down to improve cold starts and multi-tenancy. In our experiments of ElastiBench, a small fraction of microbenchmarks could not be executed successfully, e.g., because they write to the local file system. On the one hand, these benchmarks could be repeated in a different environment without significantly increasing cost and duration of the entire microbenchmark suite. On the other hand, a constrained execution environment challenges some assumptions benchmark authors make, e.g., why should the benchmark expect to be able to write to the local file system at all? Running ElastiBench could provide an interesting avenue of such feedback for microbenchmark maintainers.
8. Related Work
@@
Quickly provisioning cloud resources is a convenient approach for running software performance benchmarks on powerful infrastructure without upfront investments (Grambow et al., 2019; Javed et al., 2020). Cloud variability, however, has necessitated special approaches to ensure reliable benchmarking results (Laaber et al., 2019; Leitner and Cito, 2016; Abedi and Brecht, 2017).
@@
FaaS has proven to be useful for speeding up embarrassingly parallelizable problems in other contexts (Werner et al., 2018; Mileski and Gusev, 2022), yet to the best of our knowledge ElastiBench is the first system to run general software microbenchmarks on FaaS platforms. Previous research of benchmarking on FaaS platforms has focused on benchmarking the performance of the FaaS platform itself, e.g., in the form of functions, e.g., (Eismann et al., 2022; Lambion et al., 2022; Copik et al., 2021; Ginzburg and Freedman, 2021; Scheuner et al., 2022), or entire FaaS applications, e.g., (Grambow et al., 2021b, 2023b).
@@
Further research on improving microbenchmark execution speed, cost, and reliability has focused on changing the duration of microbenchmarks, e.g., stopping experiments early (Japke et al., 2023; He et al., 2019; Mittal et al., 2023), or running only a selection of microbenchmarks without losing information on performance regressions (Laaber et al., 2021b, 2022; Mostafa et al., 2017; De Oliveira et al., 2017; Grambow et al., 2021a). As these approaches could easily be integrated with the FaaS execution backend of ElastiBench, we consider them orthogonal to our approach.
9. Conclusion

### v2

@@ VI-A Experiment Design
Our evaluation focuses on the open-source time-series database software VictoriaMetrics, as we have used it in previous work and have reliable performance results to compare ElastiBench to [5]. We repeat the same microbenchmarks with the same SUT versions that were used in the original paper and compare their results.
@@
For this evaluation, all microbenchmarks which collect less than 10 results are ignored. Microbenchmarks might fail because they cannot run (e.g., a build failure), or because they ran for more than twenty seconds, after which they are interrupted. This can happen for some benchmarks that require extensive setups or have long iterations. We argue that these are not relevant as single microbenchmark executions microbenchmarks longer than one second with the default Go benchmark parameters are indicative of faulty benchmark implementations [13]. In an experiment, multiple microbenchmarks are run. Some microbenchmark executions are run for different configurations, e.g., different input sizes for a function. We treat these configurations as independent microbenchmarks. If the 99% confidence interval (CI) of the median difference in performance between the two SUT versions that we calculated using bootstrapping (see Section II) does not overlap 0, then the experiment has detected a performance change for that microbenchmark. If one experiment detects a performance change and another one does not, we call this a possible performance change, as it was not reliably detected across multiple experiments. Fig. 3 (note the x-axis values) shows an example of a distribution that has a performance change as opposed to no performance change.
Figure 3: When comparing the difference in performance between two versions, the confidence interval shows range where the real difference likely is. If the CI overlaps zero (black curve), we detect no performance change. If it does not overlap zero (yellow curve), we detect a performance change.
@@
Unless otherwise noted, we deploy all functions with a timeout of 15151515min, which is the maximum timeout allowed by AWS Lambda. The functions have 2048 MB RAM, and access to 1.29 vCPUs using an ARM architecture. Every function call contains one microbenchmark call, which is repeated three times inside the function. This call is repeated 15 times per microbenchmark, leading to 45 results for every microbenchmark. This is the same number used in our previous study [5] and above the number of repeats found in other research [6]. We limit the call parallelism of the call script to 150, i.e., a maximum of 150 microbenchmarks execute on the FaaS platform in parallel.
Every microbenchmark is repeated until a stable number of executions per second is found. The SUT in these experiments is the open-source time series database VictoriaMetrics. The two software versions we compare are the initial (f611434) and last commit (7ecaa2fe) evaluated in our previous work [5], which we refer to as the original dataset in this paper. In this paper, we run the following experiments:
  1. 1.
    A/A Experiment: In the A/A experiment, both versions of the SUT are the initial commit used in our previous work [5]. With this experiment, we verify that the system works and that the variability of the FaaS platform does not invalidate the results.
  2. 2.
@@ VI-B Experiment Results
We first report the results of the all experiments that collect the same data as the original dataset (Sections VI-B1, VI-B2, VI-B4, VI-B3 and VI-B5), and then compare their results against each other (Section VI-B6). Afterwards, we analyze how many repeats are necessary to achieve a consistent CI size (Section VI-B7).
VI-B1 A/A Experiment
@@
Out of 106 microbenchmarks in the SUT, 90 were executed successfully. Out of those, no performance change was found. This shows that ElastiBench does not mistakenly detect performance changes in unchanged software. Figure 4 shows an overview of the performance differences found. The median difference in performance is 0.047%, with a maximum of 32%. Large performance differences indicate that a microbenchmark is highly variable (as the same version is executed twice), which is correctly detected as no performance change. Running the experiment took ~8min and cost \$1.18555It was started ~2024-05-12T17:35:00 UTC.
@@
Figure 4: CDF showing the performance differences identified in the A/A experiment. While some microbenchmarks have a high difference, all of them are correctly categorized as no performance change.
@@
Figure 5: CDF showing the performance differences identified in the baseline experiment. Performance changes have a generally higher difference than non-changes, with the median performance change being 3.08%.
VI-B2 Baseline Experiment
@@
We show an overview of the differences in performance measured in the baseline experiment in Fig. 5. The median performance change of the two SUT versions is 4.71%. The maximum difference in performance of a change and no change are 116% and 26% respectively.
@@
After removing microbenchmarks for which only one experiment contains results, there are 91 (out of 106) microbenchmarks that can be compared to the original dataset. Out of those, both datasets agree for 95.65% of microbenchmarks (88 overall). The disagreements are as follows: for 3.2% of microbenchmarks (three overall), both find a performance change but identify different directions. These three results are from a single microbenchmark function (BenchmarkAddMulti) with different configurations. In every case, the original dataset identifies a performance change of ~-10%, while the new experiments identifies a positive 5-7% performance change. This benchmark was changed between the two SUT versions666commit hash eb103e15 with a considerable change in the benchmark itself, which leads to inconsistent results as the two versions are essentially different benchmarks. We argue that this is not representative for microbenchmarks, so that the difference in results between the original dataset and our experiments are not indicative of failure of ElastiBench.
@@
The remaining disagreement is a single microbenchmark for which the original dataset does not find a performance change while the new dataset finds one. The difference in performance is 1.96% in the original dataset, and 0.60% in the new dataset. As previous research has shown that a performance difference this small is not reliably detectable with cloud microbenchmarks (cf. Section II), we argue that while this is technically a disagreement, it is an issue of cloud benchmarking in general.
@@
The median performance change measured in this paper has a one-sided coverage with the original dataset for 86.96% of microbenchmarks, and 52.17% vice versa. The two-sided coverage is 50% of microbenchmarks, i.e., for 50% of performance changes in a microbenchmark, the result of one experiment is within the CI of the other result. We argue that the overall high agreement of the baseline experiment with the original dataset demonstrates that both find the same performance changes. The magnitude of the performance difference however depends on many factors that were changed between the baseline experiment and our original work[5], e.g., the execution environment and version of the programming language, which explains the comparatively low two-sided coverage. The experiment took ~11min, and the total cost for the experiment is ~0.18\$, i.e., a similar cost to the original experiments while taking ~4.6% of time [^7] It was started ~2024-05-12T16:50:00 UTC.
VI-B3 Replication Experiment
@@
The replication experiment has the same agreement with the original dataset as the baseline dataset. The one-sided coverage with the original dataset is 81.72%, and 51.61% in the other direction. The two-sided coverage is just 48.39% It disagrees with the baseline experiment for 10.87% of microbenchmarks, which are all microbenchmarks that are not in the original dataset (i.e., they were not run successfully or had too few runs to be included). The maximum difference in performance found by one experiment where the other experiment did not find a performance change (the possible performance change) was 5.25%. Overall, the results show that while the changes with a higher performance difference are detected continuously, smaller changes are not detected robustly. This maximum possible performance change of 5.25% is still within the limits found in other work (cf. Section II). The experiment cost \$1.18 to execute and took ~9min [^8] It was started ~2024-05-12T19:35:00 UTC.
VI-B4 Lower Memory Experiment
@@
VI-B5 Single Repeat Experiment
@@
VI-B6 Possible Performance Changes
Figure 6: The maximum difference in performance of a microbenchmark when two experiments disagree whether a performance change has happened.
We compare the difference in performance of all microbenchmarks where the baseline, replication, lower memory, and single-repeat experiments disagree with each other and collect the maximum performance change that was found by one experiment. An overview of these differences can be found in Fig. 6. The median difference is 1.58%, the 75th percentile is 3.06%, and the biggest difference is 7.6% in the BenchmarkAdd/items_100000 microbenchmark (which is unreliable, cf. Section VI-B2) between the single-repeat and lower memory experiments.
VI-B7 Repetitions Necessary for Consistent CI Size
Figure 7: If the CI of the original dataset and the experiment overlap (i.e., they show a similar performance), this graph shows how many repetitions are necessary until the size of the CI of ElastiBench is equal to or less than the size the original dataset. The graph increases until 40 repetitions and then plateaus to around 90% with more and more repetitions.
To evaluate how many repetitions of a single microbenchmark are necessary to achieve reliable results, we run an experiment where we repeat every microbenchmark 135 times: 3 times within a function call, which is called 45 times. We then calculate the median and its CI repeatedly with more and more of the calls, which leads to increasingly accurate results. Fig. 7 shows how close the results of ElastiBench can get to the original dataset. Over all benchmarks where the CIs ultimately overlap each other (i.e., they share at least one common value), we calculate how many repetitions are necessary until the size of the CI is less than or equal to the CI size of the original dataset. In 75.95% of cases, this is achieved when using the same 45 repeats used in the original dataset. At the full 135 steps, it is achieved for 89.87% of microbenchmarks.
VI-C Key Takeaways
Our evaluation of ElastiBench shows that it is viable to execute microbenchmarking experiments on FaaS platforms, decreasing microbenchmark suite execution time to 6% and lowering total execution costs compared to virtual machine-based experiments. The different experiments show that the platform and experiment configuration seem to have minimal impact on performance changes with a sufficiently large difference that they can reliably be detected using cloud platforms. For effect sizes of more than 7.06% results stayed consistent between all experiment runs, with a median inconsistent change of 1.63% (cf. Section VI-B6). Other researchers have found that a relatively small significant performance change between 3% and 10% does not imply a real-world change due to the underlying variability in cloud platforms and microbenchmarks [20, 21, 6]. Our results fall within that range.
VII Discussion & Future Work
@@
VII-A Optimizing Function Configuration
@@
We use functions with 2048 MB of memory in our evaluation to ensure no microbenchmark runs out of memory, yet not every microbenchmark necessarily requires these resources. Previous research has shown that it is possible to optimize FaaS function resource allocation further while the application is running [30]. By adapting the required resources per microbenchmark function, the total execution cost of the microbenchmark suite could be reduced further. As most cloud FaaS platforms allocate compute resources with the configured memory size, however, care must be taken to not influence benchmark results, e.g., as using shared CPU cores could further increase performance variability. Overall, however, there are limits to how low we can go in terms of resources as the SUT is, in the end, still a complete system such as VictoriaMetrics in our experiments.
VII-B Benchmarking Strategy
@@
Our current benchmarking strategy with 45 repetitions of each microbenchmark (cf. Section VI-B7) is sufficient to reduce the mean standard error of results that show a performance change to less than two percent, with an overall achievable standard error of around one percent. In future research, we plan to further refine this strategy in order to increase the cost and time savings of ElastiBench. We plan to evaluate the impact of repeating the microbenchmark within a function instance compared to repeatedly invoking a function. Additionally, we will further assess the impact of cold start performance that results from calling multiple benchmarks in parallel.
VII-C Microbenchmarks Beyond Go
@@
Our current prototype supports Go microbenchmarks, yet Go is not the only programming language with support for microbenchmarks: Future extensions to our prototype could support Rust’s cargo-bench [31], Benchmark.js [32], Java’s Microbenchmark Harness[33], or the Google Benchmark Framework for C++ projects [34]. Integrating such microbenchmark frameworks into ElastiBench will require further evaluation of their interaction with cloud FaaS performance variability.
VII-D Limitations of the Lambda Runtime
@@
As described in Section III, the execution environment of cloud FaaS platforms is typically locked down to improve cold starts and multi-tenancy. In our experiments of ElastiBench, a small fraction of microbenchmarks could not be executed successfully, e.g., because they write to the local file system. On the one hand, these benchmarks could be repeated in a different environment without significantly increasing cost and duration of the entire microbenchmark suite. On the other hand, a constrained execution environment challenges some assumptions benchmark authors make, e.g., why should the benchmark expect to be able to write to the local file system at all? Running ElastiBench could provide an interesting avenue of such feedback for microbenchmark maintainers.
VIII Related Work
@@
Quickly provisioning cloud resources is a convenient approach for running software performance benchmarks on powerful infrastructure without upfront investments [35, 36]. Cloud variability, however, has necessitated special approaches to ensure reliable benchmarking results [6, 37, 14].
@@
FaaS has proven to be useful for speeding up embarrassingly parallelizable problems in other contexts [38, 39], yet to the best of our knowledge ElastiBench is the first system to run general software microbenchmarks on FaaS platforms. Previous research of benchmarking on FaaS platforms has focused on benchmarking the performance of the FaaS platform itself, e.g., in the form of functions, e.g., [40, 41, 28, 42, 43], or entire FaaS applications, e.g., [44, 45].
@@
Further research on improving microbenchmark execution speed, cost, and reliability has focused on changing the duration of microbenchmarks, e.g., stopping experiments early [46, 47, 48], or running only a selection of microbenchmarks without losing information on performance regressions [49, 50, 51, 52, 11]. As these approaches could easily be integrated with the FaaS execution backend of ElastiBench, we consider them orthogonal to our approach.
IX Conclusion